# Synthesizing Efficient and Permissive Programmatic Runtime Shields for Neural Policies

### v1

@@ 4. Evaluation
CartPole $|\eta|<90^{\circ}\wedge|\delta|<0.3\wedge$ $|v_{1}|<0.3\wedge|v_{2}|<0.5$ 4/1 3 0.7 M
Self-Driving $|\eta|<90^{\circ}\wedge|d|<2.0$ 2/1 3 0.7 M
Pendulum $|\eta|<90^{\circ}\wedge|\omega|<\pi/2$ 2/1 3 7.0 M
4.1. Experimental Setup
Experimental Subjects
To comprehensively evaluate Aegis, we select four control systems that are heavily utilized in previous literature (Zhu et al., 2019), including autonomous driving navigation (Self-Driving), aerial drone control (Quadcopter) and dynamic equilibrium control (CartPole and Pendulum). Table 1 provides a brief description of each control system and its safety property.
The Quadcopter control system aims to realize a stable flight of a quadcopter, in which safety is ensured by limiting the angles of both the pitch and yaw rotations to not exceed $\pi/2$. The CartPole system consists of a pole attached to an unactuated joint connected to a cart moving along a frictionless track. The system is safe if the pole’s angle $\eta$ is no more than 90 from being upright, and the distance $\delta$ of the cart from the origin is less than 0.3 meters. Moreover, the velocity of the pole’s tip and cart should be less than 0.3 and 0.5 meters/s, respectively. The Pendulum system aims to balance a pendulum in an upright position. The safety is ensured by limiting the angle $\eta$ of the pendulum to not exceed 90 and the angular velocity $\omega$ to not exceed $\pi/2$. The Self-Driving system drives a car moving at a constant speed, and the safety property of this system is to ensure that the car never veers into canals found on both sides of the road, i.e, it is safe if the car’s heading angle $\eta$ remains within a threshold of $90^{\circ}$ and the distance $d$ between the car’s front and the road’s centerline stays below $2.0$ meters.
While these systems may appear simple, they are still representative of practical control systems that are difficult to control and maintain safety, such as Segway transporters (Segway, [n.d.]), camera drones (DJI, [n.d.]), and autonomous driving cars (Udacity, [n.d.]). Therefore, we believe that these control systems are appropriate for evaluating Aegis and the baseline.
We follow Zhu et al. (Zhu et al., 2019) to train the neural policies for the four control systems using the Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2016), a popular DRL algorithm for continuous control tasks. All hyperparameters are set to the default values from Zhu et al.’s official implementation. Note that our method is agnostic to the choice of DRL algorithms; other algorithms could also be applied. We choose DDPG because it is widely used in the literature (Wang et al., 2022) and aligns with the experimental setup of our baseline. We use the same neural network architecture as (Zhu et al., 2019) for each control system, which consists of 2 or 3 fully connected layers. The neural policies range in size from 0.4M to 7M neurons, as detailed in Table 1. Notably, these sizes are significantly larger than those in previous work, such as (Zhang et al., 2022; Tran et al., 2020), where policies typically contain no more than 400 neurons.
@@ 4.1. Experimental Setup
For each control system, we evaluate the performance of Aegis and the baseline VRL (Zhu et al., 2019) by synthesizing programmatic runtime shields for the neural policies. Note that VRL’s shields apply barrier certificates with high-degree polynomials while our Aegis synthesizes shields with lightweight linear policies. We follow VRL’s original implementation to set the degree of the polynomials for the barrier certificates to 4, which is the value recommended by the authors. We also tried setting the degree to 1, i.e., making the barrier certificates linear, but VRL’s synthesis process failed to converge and cannot find shields that can guarantee safety. Therefore, we use the default setting of VRL to ensure a fair comparison.
@@
Additionally, to mitigate the impact of randomness, we repeat each experiment 10 times and report the average results along with their statistical significance and effect size as suggested by (Arcuri and Briand, 2011). We use the Mann-Whitney U test (Mann and Whitney, 1947) to measure the statistical significance and the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. All the experiments are conducted on an Ubuntu 18.04 server equipped with an Intel Xeon E5-2698 CPU and 504 GB RAM.
@@ 4.2. RQ1: Effectiveness of Aegis
Pendulum 41/1000 0/1000 0/1000
Cart-Pole 47/1000 0/1000 0/1000
Self-Driving 35/1000 0/1000 0/1000
Average
76/1000
(7.6% Unsafe)
0/1000
(100% Safe)
0/1000
(100% Safe)
This RQ evaluates the effectiveness of Aegis in ensuring the safety of neural policies in the four control systems. We follow Zhu et al. (Zhu et al., 2019) to run the neural policy in each system over 1,000 runs, and each run is 5,000 time steps with random initial states. Then, we compare the number of safety violations found in the systems using neural policies, the policies with the baseline’s shields, and the policies with Aegis’s shields, as shown in Table 2.
We find that the neural policy violates the safety specifications in all the control systems, at a rate of 7.6% on average, i.e., 76 safety violations in 1,000 runs. However, after the Aegis’s and VRL’s shields are deployed, all the safety violations are eliminated, i.e., the number of violations in safety specifications is reduced to 0 in all the control systems, which suggests that both Aegis and the baseline can effectively enforce the safety by correcting all the unsafe commands and let the systems operate within the safety specifications.
@@ 4.3. RQ2: Overhead of Aegis
A programmatic runtime shield should be able to respond quickly without affecting other processes in the system, so we compare the time overhead of the shields synthesized by Aegis and VRL on the four control systems first. We measure the time overhead as the percentage of the additional operating time required by the deployed shields in the systems, i.e., if the original operating time of the system is 100s and it takes 101s after the shield is deployed, then the time overhead is 1% ((101-100)/100). Table 3 shows the time overhead of Aegis compared to the baseline. On average, the programmatic runtime shield synthesized by Aegis incurs only 5.0% time overhead, which is 5.5% lower than the 10.5% time overhead incurred by the baseline. The results show that Aegis improves the time overhead by 2.1$\times$ on average compared to the baseline.
To account for the impact of randomness and ensure the credibility of the results, we use the Mann-Whitney U test (Mann and Whitney, 1947), which is a non-parametric statistical test used to compare two independent groups. We use it to measure the statistical significance of the difference in time overhead between the runtime shields of the two tools across the 10 experiments. Our results show that Aegis significantly outperforms VRL on all four control systems at a 99% confidence level. We also use the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. This non-parametric measure gives the probability that a randomly chosen value from one group is higher or lower than one from another group. It is commonly used to evaluate search-based testing methods due to their randomness (Palomba et al., 2016; Meng et al., 2022; Olsthoorn et al., 2020). According to Vargha and Delaney (Vargha and Delaney, 2000), $\hat{A}_{12}$ values greater than 0.71 (or less than 0.29) indicate a “large” effect size. Values between 0.64 and 0.71 (or between 0.29 and 0.36) correspond to a “medium” effect size, while values between 0.36 and 0.64 suggest a “negligible” effect size. As shown in Table 2, Aegis significantly outperforms VRL with a large effect size on four control systems.
Control System Time Overhead Memory Overhead
VRL Aegis p-value Â12 VRL Aegis p-value Â12
Quadcopter 12.1% 5.3% (2.3$\times$) 5.7e-3 < 0.01 L (0.97) 1.98 MB 0.53 MB (3.7$\times$) 0.03 < 0.05 M (0.73)
Pendulum 11.9% 4.1% (2.9$\times$) 9.1e-5 < 0.01 L (1.00) 4.05 MB 0.76 MB (5.3$\times$) 2.2e-4 < 0.01 L (0.97)
Cart-Pole 10.5% 4.9% (2.1$\times$) 8.5e-4 < 0.01 L (0.87) 1.49 MB 0.58 MB (2.6$\times$) 2.9e-4 < 0.01 L (0.96)
Self-Driving 7.4% 5.5% (1.3$\times$) 7.1e-3 < 0.01 L (0.80) 3.13 MB 0.57 MB (5.5$\times$) 6.0e-3 < 0.01 L (1.00)
Average 10.5% 5.0% (2.1$\times$) 2.66 MB 0.61 MB (4.4$\times$)
@@
We then compare the memory overhead of the shields synthesized by Aegis and the baseline after deploying them into the four control systems. We measure the memory overhead as the percentage of the additional memory required by the deployed shields in the systems. Table 3 shows the memory overhead of Aegis compared to the baseline. On average, the programmatic runtime shield synthesized by Aegis occupies only 0.61 MB memory, while the baseline’s shields occupy 2.66 MB memory, which suggests that Aegis greatly reduces the memory overhead by 4.4 times compared to the baseline.
Similarly, we run each experiment 10 times and use the Mann-Whitney U test to measure the statistical significance and the Vargha-Delaney statistic $\hat{A}_{12}$ to measure the effect size of improvements. Our results show that Aegis significantly outperforms VRL on all four control systems at a 99% confidence level in terms of memory overhead, with a large or medium effect size.
RQ2 Takeaway.   The shields synthesized by Aegis significantly outperform those of the baseline in terms of time and memory overhead on all control systems. Specifically, Aegis’s shields incur only 5.0% time overhead and occupy only 0.61 MB memory on average, which are 2.1$\times$ and 4.4$\times$ improvements over the baseline, respectively.
@@ 4.4. RQ3: Permissiveness of Aegis
We first compare the permissiveness of the shields synthesized by Aegis and VRL in terms of the number of shield interventions over 1,000 runs (each run has 5,000 time steps) on the four control systems. Table 4 shows the number of shield interventions imposed by the shields of each tool. On average, Aegis requires only 384 shield interventions to ensure safety, which is 1.6$\times$ fewer than the 609 shield interventions required by the baseline. The statistical significance and effect size of the number of shield interventions are also shown in the last two columns of Table 4. Our results show that the improvements are statistically significant at the 99% confidence level with a large effect size on all four control systems.
In addition, we compare the ratio of the number of necessary shield interventions to the total number of interventions, shown in Figure 6. Here a necessary shield intervention is one that prevents a violation of the safety specifications, while an unnecessary intervention occurs when the safety specifications would not be violated even without the intervention. We observe that for each control system, Aegis’s shields impose a higher ratio of necessary shield interventions, suggesting that Aegis’s shields tend to intervene less frequently than the baseline’s shields and only when necessary. The ratio of necessary shield interventions to total number of interventions is on average 9.9% higher for Aegis than for the baseline. These results demonstrate that Aegis’s shields are more permissive than the baseline’s shields.
@@
System # Shield Interventions
@@
Quadcopter 296 232 (1.3$\times$) 8.3e-3 < 0.01 L (0.92)
Pendulum 50 45 (1.1$\times$) 8.6e-5 < 0.01 L (1.00)
Cart-Pole 1941 1133 (1.7$\times$) 9.8e-5 < 0.01 L (1.00)
Self-Driving 147 127 (1.2$\times$) 6.1e-3 < 0.01 L (1.00)
Average 609 384 (1.6$\times$)
@@
RQ3 Takeaway.  Aegis’s shields significantly outperform those of baseline in terms of permissiveness on all control systems. Specifically, Aegis requires only 384 shield interventions on average, which is a 1.6$\times$ improvement over the baseline. Furthermore, Aegis’s shields have a 9.9% higher ratio of necessary shield interventions to total number of interventions than the baseline’s shields.
@@ 4.5. RQ4: Scalability of Aegis
In the last RQ, we evaluate the scalability of Aegis by comparing the time required to synthesize a shield for a neural policy. Table 5 shows the time cost of Aegis compared to the baseline. We observe that Aegis requires 65s on average, which reduces 77.3% of the time cost compared to the baseline. The statistical significance and effect size of the time cost also show that the improvements are statistically significant at the 99% confidence level with a large effect size on all 4 control systems.
Beyond Table 5, we also draw a line chart to compare the time required to synthesize a shield for neural policies of different sizes, shown in Figure 7. As Zhu et al. (Zhu et al., 2019) provides neural policies of different sizes for the four control systems, we use these neural policies to evaluate the scalability of Aegis. We divide the neural policies into three groups based on their sizes: small (0.4M neurons), medium (0.7M neurons), and large (7M neurons). Note that such sizes are sufficiently large for neural policies, especially when contrasted with policies used in previous work such as (Zhang et al., 2022; Tran et al., 2020), which typically contain no more than 400 neurons.
In Figure 7, we can see that the synthesis time of Aegis does not significantly increase with the size of the neural policies, while the baseline’s synthesis time increases greatly when the size of the neural policies increases. This result demonstrates that Aegis’s synthesis is more scalable than the baseline’s synthesis.
System Time Cost of Shield Synthesis
VRL Aegis p-value Â12
Quadcopter 28s 14s (-50.0%) 2.2e-4 < 0.01 L (0.97)
Pendulum 439s 85s (-80.6%) 1.8e-4 < 0.01 L (1.00)
Cart-Pole 395s 118s (-70.1%) 9.3e-5 < 0.01 L (1.00)
Self-Driving 281s 43s (-84.7%) 9.5e-5 < 0.01 L (1.00)
Average 286s 65s (-77.3%)
RQ4 Takeaway  Aegis consistently shows better scalability than the baseline on all control systems. Aegis requires 65s on average for synthesizing a shield for a neural policy, which reduces 77.3% of the time cost compared to the baseline. Furthermore, Aegis’s synthesis time does not significantly increase with the size of the neural policies.

### v3

@@ 4.1. Experimental Setup
CartPole-v1 $|\delta|<0.3\;\wedge\;|\delta^{\prime}|<0.5\;\wedge$ $|\eta|<30^{\circ}\;\wedge\;|\eta^{\prime}|<0.5$ 4/1 3 0.7 M
CartPole-v2 $|\delta|<0.3\;\wedge\;|\delta^{\prime}|<0.5\;\wedge$ $|\eta|<30^{\circ}\;\wedge\;|\eta^{\prime}|<0.5$ 4/1 3 7 M
Pendulum-v1 $|\eta|<30^{\circ}\wedge|\omega|<30^{\circ}$ 2/1 3 7.0 M
Pendulum-v2 $|\eta|<30^{\circ}\wedge|\omega|<30^{\circ}$ 2/1 3 7.0 M
Self-Driving-v1 $|\eta|<90^{\circ}\wedge|d|<2.0$ 2/1 3 0.7 M
Self-Driving-v2 $|\eta|<90^{\circ}\wedge|d|<1.5$ 2/1 3 0.7 M
4-Car Platoon $|a|<=2.0\wedge|b|<=0.5\wedge$ $|c|<=0.35\;\wedge\;|d|<=0.5\;\wedge$ $|e|<=1.0\;\wedge\;|f|<=0.5\;\wedge|g|<=1.0$ 7/4 3 7.0 M
While these systems may appear simple, they are still representative of practical control systems that are difficult to control and maintain safety, such as Segway transporters (Segway, [n.d.]), camera drones (DJI, 2025), and autonomous driving cars (Udacity, 2025). Therefore, we believe that these control systems are appropriate for evaluating Aegis and the baseline.
We follow Zhu et al. (Zhu et al., 2019) to train the neural policies for the eight control systems using the Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2016), a popular DRL algorithm for continuous control tasks. All hyperparameters are set to the default values from Zhu et al.’s official implementation. Note that our method is agnostic to the choice of DRL algorithms; other algorithms could also be applied. We choose DDPG because it is widely used in the literature (Wang et al., 2024) and aligns with the experimental setup of our baseline. We use the same neural network architecture as (Zhu et al., 2019) for each control system, which consists of 2 or 3 fully connected layers. The neural policies range in size from 0.4M to 7M neurons, as detailed in Table 1. Notably, these sizes are significantly larger than those in previous work (Zhang et al., 2023; Tran et al., 2020), where policies typically contain no more than 400 neurons.
@@
For each control system, we evaluate the performance of Aegis and the baseline VRL (Zhu et al., 2019) by synthesizing programmatic runtime shields for the neural policies. Note that VRL’s shields apply barrier certificates with high-degree polynomials while our Aegis synthesizes shields with lightweight linear policies. We follow VRL’s original implementation to set the degree of the polynomials for the barrier certificates to 4, which is the value recommended by the authors. We also tried setting the degree to 1, i.e., making the barrier certificates linear, but VRL’s synthesis process failed to converge and cannot find shields that can guarantee safety. Therefore, we use the default setting of VRL to ensure a fair comparison. Moreover, the hyperparameter for inferring environment dynamics, i.e., $\epsilon$ in Algorithm 1, is set to 1.49e-10, following the default configuration used in widely-used scientific computing libraries such as SciPy (sci, 2025).
@@
Additionally, to mitigate the impact of randomness, we repeat each experiment 10 times and report the average results along with their statistical significance and effect size as suggested by Arcuri and Briand (Arcuri and Briand, 2011). We use the Mann-Whitney U test (Mann and Whitney, 1947) to measure the statistical significance and the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. All the experiments are conducted on an Ubuntu 18.04 server equipped with an Intel Xeon E5-2698 CPU and 504 GB RAM.
@@ 4.2. RQ1: Effectiveness of Aegis
Pendulum-v1 41/1000 0/1000 0/1000
Pendulum-v2 166/1000 0/1000 0/1000
Cart-Pole-v1 47/1000 0/1000 0/1000
Cart-Pole-v2 96/1000 0/1000 0/1000
Self-Driving-v1 35/1000 0/1000 0/1000
Self-Driving-v2 60/1000 0/1000 0/1000
4-Car Platoon 7/1000 0/1000 0/1000
Average 79/1000 (7.9% Unsafe) 0/1000 (100% Safe) 0/1000 (100% Safe)
This RQ evaluates the effectiveness of Aegis in ensuring the safety of neural policies across eight control systems. Following Zhu et al. (Zhu et al., 2019), we run each neural policy for 1,000 episodes, each consisting of 5,000 time steps with randomized initial states. We compare the number of safety violations under three settings: using the neural policy alone, using the baseline’s shield, and using the shield synthesized by Aegis, as summarized in Table 2.
The results show that unshielded neural policies violate safety specifications in all systems, with an average violation rate of 7.9%, corresponding to 79 violations per 1,000 runs. This highlights the inherent risks of deploying neural policies without formal safety enforcement, especially in stochastic or dynamic environments where policy behaviors may deviate from expected safe trajectories.
When either Aegis or the baseline’s shield is applied, all safety violations are eliminated across all systems—that is, the number of violations drops to 0—demonstrating that both methods effectively intercept and correct unsafe commands at runtime. The fact that Aegis achieves the same level of safety as the baseline confirms its ability to synthesize shields of equivalent effectiveness. These results also suggest that Aegis is generalizable, successfully enforcing safety across a diverse range of control systems with varying specifications and neural policy sizes, thereby underscoring its potential for deployment in real-world safety-critical applications.
@@ 4.3. RQ2: Overhead of Aegis
A programmatic runtime shield should be able to respond quickly and consume minimal resources, so we compare the time and memory overhead of the shields synthesized by Aegis and VRL across the eight control systems.
Control System Time Overhead (s) Memory Overhead (MB)
VRL Aegis p-value Â12 VRL Aegis p-value Â12
[…]
@@
[…]
To ensure the robustness of our findings and account for randomness, we apply the Mann-Whitney U test (Mann and Whitney, 1947), a non-parametric statistical test for comparing two independent groups. We use this test to evaluate whether the difference in time overhead between the two methods is statistically significant across 10 experimental runs. As shown in Table 3, the time overhead of Aegis is statistically significantly lower than that of VRL, with all p-values well below 0.01, indicating statistical significance at the 99% confidence level and a very low probability (less than 1%) of being due to random chance.
We also report the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to quantify the effect size. This non-parametric measure gives the probability that a randomly chosen value from one group is higher or lower than one from another group. It is commonly used to evaluate search-based testing methods due to their randomness (Meng et al., 2022; Olsthoorn et al., 2021). According to Vargha and Delaney (Vargha and Delaney, 2000), $\hat{A}_{12}$ values greater than 0.71 (or less than 0.29) indicate a “large” effect size. Values between 0.64 and 0.71 (or between 0.29 and 0.36) correspond to a “medium” effect size, while values between 0.36 and 0.64 suggest a “negligible” effect size. As shown in Table 3, Aegis outperforms VRL with a large effect size across all eight control systems, with all $\hat{A}_{12}$ values exceeding 0.80. This reflects a strong and consistent advantage of Aegis over VRL– in at least 80% of the experiments, Aegis incurs lower time overhead than VRL.
@@
[…]
Similarly, we run each experiment 10 times and use the Mann-Whitney U test to validate the statistical significance and the Vargha-Delaney statistic $\hat{A}_{12}$ to measure the effect size of improvements. Our results show that Aegis significantly outperforms VRL on all eight control systems at a 99% confidence level in terms of memory overhead, except for the Quadcopter system, where the improvement is still statistically significant at the 95% confidence level. The effect size of the memory overhead is also large for all control systems, with $\hat{A}_{12}$ values exceeding 0.73, indicating that Aegis’s shields are consistently more efficient than those of the baseline across all systems with low probability of being due to random chance.
[…]
@@ 4.4. RQ3: Permissiveness of Aegis
Control System # Shield Interventions
@@
[…]
Similar to the analysis in RQ2, we have conducted 10 independent runs and applied the Mann-Whitney U test to assess statistical significance. As shown in the last two columns of Table 4, all eight systems show $p$-values below 0.01, indicating the difference is unlikely due to chance. Furthermore, the Vargha-Delaney $\hat{A}_{12}$ effect sizes are large in most systems, reaching 1.00 in five cases, confirming that Aegis’s improvement is both statistically and practically significant.
We also compare the ratio of necessary shield interventions to total interventions, as shown in Figure 12. A necessary intervention is one that prevents a violation of the safety specifications, while an unnecessary one occurs when no violation would have happened even without intervention. Notably, the number of necessary interventions does not always equal the number of safety violations observed under the original (unshielded) neural policy, as reported in RQ1. This discrepancy arises because, once a shield intervenes, it modifies the system’s trajectory by altering the original control command. While this new trajectory may avoid some original violations, it can also introduce new potential violations at other time steps that would not have occurred without shielding. Across all systems, Aegis’s shields consistently impose a higher ratio of necessary interventions, averaging 10.2% more than the baseline. This indicates that Aegis’s shields are more precise, intervening only when safety is genuinely at risk, rather than interrupting benign commands from neural policies. Such precision is crucial for maintaining system stability while ensuring safety. A permissive shield improves overall stability, further reduces actuation delays, and increases trust in the policy during real-world deployment.
QuadcopterPendulum-v1Pendulum-v2Cart-Pole-v1Cart-Pole-v2Self-Driving-v1Self-Driving-v24-Car Platoon0%25%50%75%100%66.6%82.0%71.4%38.5%50.3%21.8%39.0%77.8%75.1%95.6%89.1%46.9%55.1%30.7%48.0%88.2%VRL (avg. 55.9%) Aegis (avg. 66.1%)
@@
[…]
Control System Time Cost of Shield Synthesis
VRL Aegis p-value Â12
[…]
Small (0.4 M)Medium (0.7 M)Large (7 M)0s125s250s375s500sAegisVRLScale of Neural NetworkSynthesis Time (s)
@@ 4.5. RQ4: Scalability of Aegis
[…]
To check for statistical significance, we perform 10 independent runs per system and apply the Mann-Whitney U test. All $p$-values are well below 0.01, confirming the improvements are statistically significant. The Vargha-Delaney $\hat{A}_{12}$ values also indicate large effect sizes across all systems, showing that the improvements are substantial in practice.
Beyond Table 5, we also draw a line chart to compare the time required to synthesize a shield for neural policies of different sizes, shown in Figure 13. As Zhu et al. (Zhu et al., 2019) and Shi et al. (Shi et al., 2025b) provide neural policies of different sizes for the eight control systems, we use these neural policies to evaluate the scalability of Aegis. We divide the neural policies into three groups based on their sizes: small (0.4M neurons), medium (0.7M neurons), and large (7M neurons). Note that such sizes are sufficiently large for neural policies, especially when contrasted with policies used in previous work such as (Zhang et al., 2023; Tran et al., 2020), which typically contain no more than 400 neurons.
Figure 13 shows that Aegis’s synthesis time remains relatively stable as the size of the neural policies increases, whereas the baseline’s synthesis time grows significantly. This result demonstrates that Aegis scales more effectively with policy size and handles increasing complexity gracefully. In contrast, the baseline’s time increases sharply, likely due to the greater burden of constraint solving in more complex settings. This scalability is crucial in practice, as modern applications in robotics, autonomous driving, and cyber-physical systems increasingly rely on large neural policies. Aegis’s ability to handle such policies without a substantial increase in synthesis time makes it a more practical solution for real-world deployment.
[…]